OpenStudy (anonymous):

You kick a ball with a speed of 4 m/s at an angle of 10 degrees. How far away does the ball land?

OpenStudy (anonymous):

in Y Vyfinal= Vyi-9.81 ty 0= (4sin10)^2-2*9.8 ty to reach the highest point solve ty in x Vx=x/2ty 4cos10=x/2ty solve x